OpenStudy (adamconner):

Find the point C guaranteed by the mean value theorem for f(x)? 1 MEDAL & FAN f(x)=(x+3)/x on the interval [1,6]. so a=1 and b=6 f(1)=4 and f(6)=3/2 after this part i don't know what to do.

OpenStudy (anonymous):

You know what mean value theorem states?

OpenStudy (anonymous):

If you don't, then first check it. if you do, then find the slope of the cord between (1,4) and (6,3/2) then find derivative of f(x), set it equal to the slope you found previously and solve for x. That's it

OpenStudy (adamconner):

so -1/2=3/x^2??

OpenStudy (anonymous):

slope=\(\large\frac{3/2-4}{6-1}=-1/2\) derivative=\(\large\frac{-3}{x^2}\) so: \(3/x^2=1/2\) or x^2=6 so \(x=\pm \sqrt{6}\)

OpenStudy (phi):

of course you only want the positive root, because you want x on the interval [1,6].

OpenStudy (adamconner):

thanks guys